# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 5 Anytime-valid, Bayes-assisted, prediction-powered inference
In this section, we show how the results of Sections 3 and 4 can be combined in the context of PPI to obtain AsympCS for $g_{\theta}$. For any $\theta\in\mathbb{R}$ and $i\geq 1$, let $U_{\theta,i}=\ell_{\theta}^{\prime}(X_{i},f(X_{i}))$, $\widetilde{U}_{\theta,i}=\ell_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i}))$ and $V_{\theta,i}=\ell_{\theta}^{\prime}(X_{i},Y_{i})$. Define $\overline{V}_{\theta,n}=\frac{1}{n}\sum_{i=1}^{n}V_{\theta,i}$ and $\overline{U}_{\theta,n}=\frac{1}{n}\sum_{i=1}^{n}U_{\theta,i}$.
@@ 5.1 Anytime-valid PPI
We first derive AsympCS that do not incorporate prior information on the accuracy of the black-box predictor. The following proposition follows directly from Proposition 4.2 and Theorem 3.4, owing to the control-variate form of the PPI estimator $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$ (17) and PPI++ estimator $\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}$ (20).
Proposition 5.1.
Assume $|n/N_{n}-r|=O(1/\log n)$ for some $r\in[0,1]$. Let $\widehat{g}_{\theta,n}$ be either the PPI (17) or the PPI++ (20) estimators. For PPI, let $(\widehat{\sigma}^{g}_{\theta,n})^{2}=\widehat{\nu}_{1}^{\mathrm{cv}}((U_{\theta,i},V_{\theta,i})_{i=1}^{n},(\widetilde{U}_{\theta,i})_{i=1}^{N_{n}})$ (see (13)). For PPI++, let $(\widehat{\sigma}^{g}_{\theta,n})^{2}=\widehat{\nu}^{\mathrm{cv+}}((U_{\theta,i},V_{\theta,i})_{i=1}^{n})$ (see (14)). Then, for any parameter $\rho$, the sequence of intervals defined as $\mathcal{C}^{g}_{\alpha,\theta,n}=\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\widehat{g}_{\theta,n},\widehat{\sigma}^{g}_{\theta,n};\rho)$ forms a $(1-\alpha)$–AsympCS for $g_{\theta}$.
@@ 5.2 Anytime-valid, Bayes-assisted, PPI
In many modern applications, extremely accurate black-box predictors are available. When this is the case, we can then leverage this prior information to obtain tighter AsympCS for $g_{\theta}$ via a zero-mean prior on $\Delta_{\theta}$. Following the decomposition in Equation 3, we simply combine an AsympCS for $m_{\theta}$ (Proposition 5.2) with a Bayes-assisted AsympCS for $\Delta_{\theta}$ (Proposition 5.3).
Proposition 5.2 (AsympCS for $m_{\theta}$).
@@
Let $\widehat{m}_{\theta,n}$ be the sample mean estimator (15) and $(\widehat{\sigma}^{f}_{\theta,n})^{2}$ be the sample variance of $(\ell_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i})))_{i=1}^{N_{n}}$. Let $\delta\in(0,1)$. For any $\rho>0$, $\mathcal{R}_{\delta,\theta,n}=\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\delta,n}(\widehat{m}_{\theta,n},\widehat{\sigma}^{f}_{\theta,n};\rho)$ forms a $(1-\delta)$–AsympCS for $m_{\theta}$.
Proposition 5.3 (Bayes-assisted AsympCS for $\Delta_{\theta}$).
@@
Let $\widehat{\Delta}_{\theta,n}$ be either the sample mean estimator (16) (PPI) or the control variate estimator (19) (PPI++). For PPI, let $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ be the sample variance of $(V_{\theta,i}-U_{\theta,i})_{i=1}^{n}$. For PPI++, let $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}=\widehat{\nu}^{\mathrm{cv+}}((U_{\theta,i},V_{\theta,i}-U_{\theta,i})_{i=1}^{n})$ (see (14)). Let $\kappa\in(0,1)$ and assume that $|\frac{n}{N_{n}}-r|=O(1/\log n)$ for some $r\in[0,1]$. For any continuous proper prior $\pi$, the sequence of Bayes-assisted intervals $\mathcal{T}_{\kappa,\theta,n}=\mathcal{C}_{\kappa,n}^{\mathrm{\mathtt{B}A}}(\widehat{\Delta}_{\theta,n},\widehat{\sigma}^{\Delta}_{\theta,n};\pi)$ forms a $(1-\kappa)$–AsympCS for $\Delta_{\theta}$.
@@
Finally, for both PPI and PPI++, the confidence sequences $\mathcal{R}_{\delta,\theta,n}$ and $\mathcal{T}_{\alpha-\delta,\theta,n}$ are combined via a Minkowski sum to obtain a $(1-\alpha)$–AsympCS for $g_{\theta}$, of the form
@@
$\mathcal{C}_{\alpha,\theta,n}^{g}=\left[\widehat{g}_{\theta,n}\pm\left\{\frac{\widehat{\sigma}^{\Delta}_{\theta,n}}{\sqrt{n}}\sqrt{\log\left(\frac{n(2\pi\kappa^{2})^{-1}}{\eta_{n}(\widehat{\Delta}_{\theta,n}/\widehat{\sigma}^{\Delta}_{\theta,n})}\right)}+\frac{\widehat{\sigma}^{f}_{\theta,n}}{\sqrt{N_{n}}}\sqrt{\frac{1+N_{n}\rho^{2}}{N_{n}\rho^{2}}\log\left(\frac{N_{n}\rho^{2}+1}{\delta^{2}}\right)}\right\}\right]$ (22)
@@
where $\widehat{g}_{\theta,n}$ is either the PPI estimator (17) or the PPI++ estimator (20). Solving Equation 4 gives the confidence region for $\theta^{*}$. In the case of the squared loss, $\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ is an interval, given by
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}=\left[\widehat{\theta}_{n}\pm\left\{\frac{\widehat{\sigma}^{\Delta}_{0,n}}{\sqrt{n}}\sqrt{\log\left(\frac{n}{2\pi\kappa^{2}\eta_{n}(\widehat{\Delta}_{0,n}/\widehat{\sigma}^{\Delta}_{0,n})}\right)}+\frac{\widehat{\sigma}^{f}_{\theta,n}}{\sqrt{N_{n}}}\sqrt{\frac{1+N_{n}\rho^{2}}{N_{n}\rho^{2}}\log\left(\frac{N_{n}\rho^{2}+1}{\delta^{2}}\right)}\right\}\right]$ (23)
@@ 6 Experiments
We compare the PPI and PPI++ AsympCS procedures introduced in Section 5 – with and without Bayes-assistance – to the AsympCS relying solely on labelled data (obtained from Theorem 3.4 and referred to as “classical”) on scalar mean estimation problems, i.e. $\theta^{*}=\mathbb{E}[Y]\in\mathbb{R}$. Bayes-assisted methods are annotated as (G), (L), or (T) to indicate Gaussian, Laplace, or Student-t priors with mean zero and scale depending on the task and reported in the supplementary material. Since PPI is motivated by settings with scarce labelled data and abundant unlabelled data, we consider the following experimental setting: labelled data arrives sequentially, i.e. $n=1,2,\ldots$, while a large unlabelled dataset is available from the start, i.e. $N_{n}=N$ for all $n$, with $N\gg n$ large enough to exclude any uncertainty on the measure of fit $m_{\theta}$. As discussed by Cortinovis and Caron (2025), this simplifies the comparison between unassisted and Bayes-assisted PPI, as it rules out any potential loss of efficiency due to the Minkowski sum (22), thereby isolating the effect of the Bayes correction on the CS procedure. For synthetic data, we set $N=\infty$ to guarantee the simplification holds. For real data, we empirically verify that $N$ is large enough to justify this assumption by confirming that anytime validity is preserved – specifically, that the cumulative miscoverage rate remains below the chosen threshold $\alpha=0.1$ for all $n$. As with CLT-based CIs, the $n$ at which one starts counting the cumulative miscoverage rate of an asymptotic CS is inherently arbitrary. Here, we choose $n=40$, as we empirically find this to be a reasonably small labelled sample size at which Strassen’s strong coupling generally provides a good approximation.
@@ 6.1 Synthetic data
The synthetic data experiments below follow a general structure: we start with $N=\infty$ unlabelled samples $\{\widetilde{X}_{j}\}_{j=1}^{N}$ i.i.d. from $\mathbb{P}_{X}$ and successively sample $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$ i.i.d. from $\mathbb{P}$. We compare methods in terms of the average interval volume, and report their cumulative miscoverage rate in the supplementary material.
@@
This experiment demonstrates that our method can adapt to varying correlation levels between predictions and true labels by using the PPI++ estimator (19). We sample $Y_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$, so that $\theta^{*}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=Y_{i}+\epsilon_{i}$, where $X_{i}$ is only used for indexing and $\epsilon_{i}\overset{iid}{\sim}\mathcal{N}(0,\sigma^{2}_{Y})$, with the noise level $\sigma_{Y}$ successively set to $0.1,0.8,3$. In this case, the optimal control-variate parameter is given by $\lambda^{*}_{\theta}=\lambda^{*}=\mathrm{cov}(Y,f(X))/\mathrm{var}(f(X))=(1+\sigma_{Y}^{2})^{-1}$, which decreases with $\sigma_{Y}$. Figure 1 compares the volume achieved by classical and unassisted CS procedures as a function of $n$, while results under informative priors are reported in the supplementary material.
Figure 1: Noisy predictions study. The left, middle and right panels show average interval volume over $1000$ repetitions for noise levels $\sigma_{Y}=0.1,0.8,3.0$.
For small noise levels, PPI and PPI++ achieve a similar result, and greatly outperform classical inference. As the noise level grows, the machine learning predictions become less informative and standard PPI loses ground to the classical CS. On the other hand, PPI++ is able to adapt to the noise level and always performs similarly or better than the other baselines.
@@
This experiment illustrates the potential benefits of incorporating prior information into our method. We sample $X_{i}\overset{iid}{\sim}\mathcal{N}(0,1)$ and $Y_{i}=X_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is sampled i.i.d. from a Student-t distribution with $\mathrm{df}$ degrees of freedom and scale $10$, so that $\theta^{*}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=X_{i}+\upsilon$, where $\upsilon\in\mathbb{R}$ controls the bias level. For all $\upsilon$, $\lambda^{*}=1$, so that PPI and PPI++ coincide. We vary $\upsilon$ between $-5$ and $5$, and $\mathrm{df}\in\{5,10,\infty\}$ to study the impact of bias level and noise distribution on the AsympCS procedures. Figure 2 compares the average interval volumes at $n=100$.
Figure 2: Biased predictions study. The left, middle and right panels show average interval volume over $100$ repetitions for $\mathrm{df}=5,10,\infty$.
Classical inference and unassisted PPI volumes remain constant across bias levels, reflecting their lack of prior information, and achieve similar results, as the estimated variances of $Y$ and $\Delta_{\theta}$ are very close due to the high scale of the noise $\epsilon$. On the other hand, the volume of the Bayes-assisted procedures varies widely with the bias level $\upsilon$: the volume is reduced for small $\upsilon$, but grows with $|\upsilon|$ as the priors become increasingly mispecified. Notably, the volume of the Gaussian prior increases the fastest with $|\upsilon|$, while heavier-tailed Laplace and Student-t priors offer comparatively greater robustness. These conclusions hold for all values of $\mathrm{df}$, which controls the accuracy of Strassen’s coupling approximation for a given $n$. Coverage results, reported in the supplementary material, show that, while smaller values of $\mathrm{df}$ lead to slightly worse coverage, the approximation quality is overall satisfactory in this example.

### v2

@@ 5 Anytime-valid, Bayes-assisted, prediction-powered inference
In this section, we show how the results of Sections˜3 and 4 can be combined in the context of PPI to obtain AsympCS for $g_{\theta}$. For any $\theta\in\mathbb{R}$ and $i\geq 1$, let $U_{\theta,i}=\ell_{\theta}^{\prime}(X_{i},f(X_{i}))$, $\widetilde{U}_{\theta,i}=\ell_{\theta}^{\prime}(\widetilde{X}_{i},f(\widetilde{X}_{i}))$ and $V_{\theta,i}=\ell_{\theta}^{\prime}(X_{i},Y_{i})$. Define $\overline{V}_{\theta,n}=\frac{1}{n}\sum_{i=1}^{n}V_{\theta,i}$ and $\overline{U}_{\theta,n}=\frac{1}{n}\sum_{i=1}^{n}U_{\theta,i}$. In the following, we assume $\mathbb{E}|U_{\theta,i}|^{2+\delta}$, $\mathbb{E}|\widetilde{U}_{\theta,i}|^{2+\delta}$ and $\mathbb{E}|V_{\theta,i}|^{2+\delta}<\infty$ for some $0<\delta<1$, and that $|\frac{n}{N_{n}}-r|=O(1/n^{1-a})$ with $0<a<2/(2+\delta)$ for some $r\in[0,1]$.
@@ 5.1 Anytime-valid PPI
We first derive AsympCS that do not incorporate prior information on the accuracy of the black-box predictor. The following result follows directly from Proposition˜2 and Theorem˜2, owing to the control-variate form of the PPI estimator $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$ (21) and the PPI++ estimator $\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}$ (24).
Proposition 3.
Let $\widehat{g}_{\theta,n}$ be either the PPI (21) or the PPI++ (24) estimator. For PPI, let $(\widehat{\sigma}^{g}_{\theta,n})^{2}=\widehat{\nu}_{1}^{\mathrm{cv}}((U_{\theta,i},V_{\theta,i})_{i=1}^{n},(\widetilde{U}_{\theta,j})_{j=1}^{N_{n}})$ (see (17)). For PPI++, let $(\widehat{\sigma}^{g}_{\theta,n})^{2}=\widehat{\nu}^{\mathrm{cv+}}((U_{\theta,i},V_{\theta,i})_{i=1}^{n})$ (see (18)). Then, for any $\rho>0$, the sequence of intervals defined as $\mathcal{C}^{g}_{\alpha,\theta,n}=\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,n}(\widehat{g}_{\theta,n},\widehat{\sigma}^{g}_{\theta,n};\rho)$ forms a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{n\log{n}}$ for $g_{\theta}$ and asymptotic Type-I error control.
@@ 5.2 Anytime-valid, Bayes-assisted, PPI
In many modern applications, extremely accurate black-box predictors are available (e.g., [26, 27, 28]). When this is the case, we can leverage this prior information to obtain tighter AsympCS for $g_{\theta}$ via a zero-mean prior on $\Delta_{\theta}$. Following the decomposition in Equation˜3, we combine an AsympCS for $m_{\theta}$ (Proposition˜4) with a Bayes-assisted AsympCS for $\Delta_{\theta}$ (Proposition˜5).
Proposition 4 (AsympCS for $m_{\theta}$).
@@
Let $\widehat{m}_{\theta,n}$ and $(\widehat{\sigma}^{f}_{\theta,n})^{2}$ be the sample mean (19) and sample variance of $(\ell_{\theta}^{\prime}(\widetilde{X}_{j},f(\widetilde{X}_{j})))_{j=1}^{N_{n}}$. Let $\delta\in(0,1)$. For any $\rho>0$, $\mathcal{R}_{\delta,\theta,n}=\mathcal{C}^{\mathrm{\tt NA}}_{\delta,n}(\widehat{m}_{\theta,n},\widehat{\sigma}^{f}_{\theta,n};\rho)$ forms a $(1-\delta)$–AsympCS with approximation rate $1/\sqrt{n\log{n}}$ for $m_{\theta}$ and asymptotic Type-I error control.
Proposition 5 (Bayes-assisted AsympCS for $\Delta_{\theta}$).
@@
For PPI, let $\widehat{\Delta}_{\theta,n}$ and $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ be the sample mean (20) and sample variance of $(V_{\theta,i}-U_{\theta,i})_{i=1}^{n}$. For PPI++, let $\widehat{\Delta}_{\theta,n}$ be the control-variate estimator (23) and $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}=\widehat{\nu}^{\mathrm{cv+}}((U_{\theta,i},V_{\theta,i}-U_{\theta,i})_{i=1}^{n})$ (see (18)). Let $\kappa\in(0,1)$. For any continuous proper prior $\uppi$, the sequence of Bayes-assisted intervals $\mathcal{T}_{\kappa,\theta,n}=\mathcal{C}_{\kappa,n}^{\mathrm{\tt BA}}(\widehat{\Delta}_{\theta,n},\widehat{\sigma}^{\Delta}_{\theta,n};\uppi)$ forms a $(1-\kappa)$–AsympCS with approximation rate $1/\sqrt{n\log{n}}$ for $\Delta_{\theta}$ and asymptotic Type-I error control.
@@
Finally, for both PPI and PPI++, the confidence sequences $\mathcal{R}_{\delta,\theta,n}$ and $\mathcal{T}_{\alpha-\delta,\theta,n}$ are combined via a Minkowski sum to obtain a $(1-\alpha)$–AsympCS for $g_{\theta}$ with approximation rate $1/\sqrt{n\log{n}}$ for $\mu$ and asymptotic Type-I error control, of the form
@@
$\mathcal{C}_{\alpha,\theta,n}^{g}=\left[\widehat{g}_{\theta,n}\pm\left\{\frac{\widehat{\sigma}^{\Delta}_{\theta,n}}{\sqrt{n}}\sqrt{\log\left(\frac{n(2\pi\kappa^{2})^{-1}}{\eta_{n}(\widehat{\Delta}_{\theta,n}/\widehat{\sigma}^{\Delta}_{\theta,n})^{2}}\right)}+\frac{\widehat{\sigma}^{f}_{\theta,n}}{\sqrt{N_{n}}}\sqrt{\frac{1+N_{n}\rho^{2}}{N_{n}\rho^{2}}\log\left(\frac{N_{n}\rho^{2}+1}{\delta^{2}}\right)}\right\}\right]$ (26)
@@
where $\widehat{g}_{\theta,n}$ is either the PPI estimator (21) or the PPI++ estimator (24). Solving Equation˜4 gives the confidence region for $\theta^{\star}$. In the case of the squared loss, $\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ is an interval, given by
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}=\left[\widehat{\theta}_{n}\pm\left\{\frac{\widehat{\sigma}^{\Delta}_{0,n}}{\sqrt{n}}\sqrt{\log\left(\frac{n}{2\pi\kappa^{2}\eta_{n}(\widehat{\Delta}_{0,n}/\widehat{\sigma}^{\Delta}_{0,n})^{2}}\right)}+\frac{\widehat{\sigma}^{f}_{0,n}}{\sqrt{N_{n}}}\sqrt{\frac{1+N_{n}\rho^{2}}{N_{n}\rho^{2}}\log\left(\frac{N_{n}\rho^{2}+1}{\delta^{2}}\right)}\right\}\right]$ (27)
@@ 6 Experiments
We compare the PPI and PPI++ AsympCS procedures introduced in Section˜5 – with and without Bayes assistance – to the AsympCS relying solely on labelled data (obtained from Theorem˜1 and referred to as “classical”) on several estimation problems. Bayes-assisted methods are annotated with (G), (L), or (T) to indicate Gaussian, Laplace, or Student-t priors with mean zero and scale depending on the task and reported in the Supplementary Material. For the Student-t prior, we set the degrees of freedom to $2$ in all experiments. Since PPI is motivated by settings with scarce labelled data and abundant unlabelled data, we consider the following experimental setting: labelled data arrive sequentially, i.e., $n=1,2,\ldots$, while a large unlabelled dataset is available from the start, i.e., $N_{n}=N$ for all $n$, with $N\gg n$ large enough to exclude any uncertainty on the measure of fit $m_{\theta}$. As discussed by Cortinovis and Caron [6], this simplifies the comparison between non-assisted and Bayes-assisted PPI, as it rules out any potential loss of efficiency due to the Minkowski sum (26), thereby isolating the effect of the Bayes correction on the CS procedure. For synthetic data, we set $N=\infty$ to guarantee the simplification holds. For real data, we empirically verify that $N$ is large enough to justify this assumption by confirming that anytime validity is preserved – specifically, that the cumulative miscoverage rate remains below the chosen threshold $\alpha=0.1$ for all $n$. As with CLT-based CIs, the $n$ at which one starts counting the cumulative miscoverage rate of an asymptotic CS is inherently arbitrary; unless otherwise stated, we choose $n=40$, as we empirically find this to be a reasonably small labelled sample size at which the KMT coupling generally provides a good approximation.
@@ 6.1 Synthetic data
The synthetic experiments below follow a general structure: we start with $N=\infty$ unlabelled samples $\{\widetilde{X}_{j}\}_{j=1}^{N}\overset{\mathrm{iid}}{\sim}\mathbb{P}_{X}$ and successively sample $n$ labelled observations $\{(X_{i},Y_{i})\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\mathbb{P}$ with the goal of estimating the mean $\theta^{\star}=\mathbb{E}[Y]$. We compare methods in terms of the average interval volume as $n$ increases across repetitions, and report the associated cumulative miscoverage rate in Section˜S7.1.
@@
This experiment demonstrates that our method can adapt to varying correlation levels between predictions and true labels by using the PPI++ estimator (23). We sample $Y_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$, so that $\theta^{\star}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=Y_{i}+\epsilon_{i}$, where $X_{i}$ is only used for indexing and $\epsilon_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\sigma_{Y}^{2})$, with the noise level $\sigma_{Y}\in\{0.1,0.8,3\}$. In this case, the optimal control-variate parameter is given by $\lambda^{\star}_{\theta}=\lambda^{\star}=\mathrm{cov}(Y,f(X))/\mathrm{var}(f(X))=(1+\sigma_{Y}^{2})^{-1}$, which decreases with $\sigma_{Y}$. Figure˜1 compares the interval volume achieved by classical and non-assisted CS procedures as a function of $n$, while results under informative priors are reported in Section˜S7.1.
Figure 1: Noisy predictions study. The left, middle and right panels show average interval volume over $1000$ repetitions as a function of the labelled sample size $n$ for noise levels $\sigma_{Y}\in\{0.1,0.8,3.0\}$.
For small noise levels, PPI and PPI++ achieve similar performance, and greatly outperform classical inference. As the noise level grows, the machine learning predictions become less informative and standard PPI loses ground to the classical CS. By contrast, PPI++ adapts to the noise level and always performs similarly to, or better than, the other baselines.
@@
This experiment illustrates the potential benefits of incorporating prior information into our method. We sample $X_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1)$ and $Y_{i}=X_{i}+\epsilon_{i}$, where $\epsilon_{i}\overset{\mathrm{iid}}{\sim}t_{\mathrm{df}}(0,1)$, so that $\theta^{\star}=\mathbb{E}[Y]=0$. The prediction rule is defined as $f(X_{i})=X_{i}+\upsilon$, where $\upsilon\in\mathbb{R}$ controls its bias level. For all $\upsilon$, $\lambda^{\star}=1$, so that PPI and PPI++ coincide. We vary $\upsilon$ between 1.2-1.2 and $1.2$, and $\mathrm{df}\in\{5,10,\infty\}$ to study the impact of bias level and noise distribution on the AsympCS procedures. Figure˜2 compares the average interval volumes at $n=100$ as a function of $\upsilon$ for each value of $\mathrm{df}$.
Figure 2: Biased predictions study. The left, middle and right panels show average interval volume over $100$ repetitions as a function of the bias level $\upsilon$ for $\mathrm{df}=5,10,\infty$.
Classical inference and non-assisted PPI volumes remain essentially constant across bias levels, reflecting their lack of prior information, and with the latter consistently outperforming the former by leveraging imputed predictions. On the other hand, the volume of the Bayes-assisted procedures varies widely with the bias level $\upsilon$: the volume is reduced for small $\upsilon$, but grows with $|\upsilon|$ as the priors become increasingly misspecified. Notably, the volume under the Gaussian prior inflates the fastest with $|\upsilon|$, while heavier-tailed Laplace and Student-t priors offer comparatively greater robustness. These conclusions hold for all values of $\mathrm{df}$, which controls the accuracy of the KMT coupling approximation for a given $n$. Coverage results in Section˜S7.1 show that, while smaller values of $\mathrm{df}$ lead to slightly worse coverage, the approximation quality is overall satisfactory in this example.